A body falls freely from rest. Find;

a. distance it falls in 5 sec.

b. velocity after falling 5km.

c. time required to attain a velocity of 50m/s

d. time in which it falls 200m.


1
Expert's answer
2021-02-25T10:53:11-0500

a)


"h=0.5gt^2=0.5(9.8)(5)^2=122.5\ m"

b)


"v=\sqrt{2gh}=\sqrt{2(9.8)(5000)}=313\frac{m}{s}"

c)


"t=\frac{v}{g}=\frac{50}{9.8}=5.1\ s"

d)


"T=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(200)}{9.8}}=6.4\ s"